OpenStudy (anonymous):

if f(x) = 5x^3 - 2 and g(x) = x + 1 find (f-g) (x) 5x^4 - 2 5x^3 -x -3 5x^3 - x -1 5x^3 + 1

OpenStudy (anonymous):

Wouldn't (f-g)(x) = (5x^3-2)-(x+1)

OpenStudy (michele_laino):

please note that: \[(f-g)(x)=f(x)-g(x)=...\] so?

OpenStudy (anonymous):

\[(f-g) (x) \] is simply \[f(x) - g(x)\]

OpenStudy (anonymous):

So, it would be (f-g)(x)=5x^3-2-x+1

OpenStudy (michele_laino):

please, be careful to signs! @GodGirl360

OpenStudy (anonymous):

Then my first equation that I set up would be correct, right @Michele_Laino. The next step would be to distribute the negative sign. (f-g)(x)=5x^3-2-x-1 Then you would combine like terms. (f-g)(x)=5x^3-x-3

OpenStudy (anonymous):

So, B is the answer?

OpenStudy (michele_laino):

please note that in your formula, namely: (f-g)(x)=5x^3-2-x+1 there is a sign error, do you agree? @GodGirl360

OpenStudy (anonymous):

Yes, that is why I corrected myself and went back to solving my initial equation, @Michele_Laino, which was (f-g)(x)=(5x^3-2)-(x+1). The rest of the work is mentioned in one of my last comments.

Nnesha (nnesha):

@Tallan did you understand ????

OpenStudy (anonymous):

So, would B be to answer?

OpenStudy (anonymous):

no im still a little confused on how you get it. So when you set the problem up (5x^3 - 2) - (x + 1) right? then you would distribute?

Nnesha (nnesha):

yes

Nnesha (nnesha):

let the asker solve this question **

OpenStudy (anonymous):

Sorry @Nnesha, got ahead of myself.

Nnesha (nnesha):

please don't be sorry its okay

OpenStudy (anonymous):

ok so my question is which number do I use the 5x^3 to distribute?

Nnesha (nnesha):

nope like she already mention above first post you have to subtract distribute bracket by -1